# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v3

@@ 3 $+T(a_{1},c_{1})$$\{(a_{1},b_{1},c_{1})\}$$\{+Q(a_{1},b_{1},c_{1})\}$
Given a query $Q$, let $\mathsf{w}(Q)$ denote the fractional hypertree width of $Q$ (see Section 4). Let $\mathsf{\omega}(Q)$ denote the smallest exponent $\alpha$ such that $\text{Eval}[Q]$ has an algorithm with runtime $\mathcal{O}(|\mathcal{D}|^{\alpha+o(1)}+|Q(\mathcal{D})|)$ on any input database $\mathcal{D}$. Given a union of queries $Q=\bigvee_{i}Q_{i}$, we define $\mathsf{w}(Q)$ and $\mathsf{\omega}(Q)$ to be the maximum of $\mathsf{w}(Q_{i})$ and $\mathsf{\omega}(Q_{i})$, respectively, over all $i$.  [^1] The function $\omega(Q)$ is defined analogously to the constant $\omega$ which is the best exponent for matrix multiplication.
The function $\mathsf{\omega}(Q)$ is not known in general. However, we know that $\mathsf{\omega}(Q)\leq\mathsf{w}(Q)$; see e.g. (Abo Khamis et al., 2016). More strongly, $\mathsf{\omega}(Q)$ is upper bounded by the submodular width of $Q$ (Marx, 2013; Abo Khamis et al., 2017; Abo Khamis et al., 2024)222 (Abo Khamis et al., 2017; Abo Khamis et al., 2024) provide a query evaluation algorithm with runtime $\mathcal{O}(|\mathcal{D}|^{\mathrm{subw}(Q)}\cdot\text{\sf polylog}{|\mathcal{D}|}+|Q(\mathcal{D})|)=\mathcal{O}(|\mathcal{D}|^{\mathrm{subw}(Q)+o(1)}+|Q(\mathcal{D})|)$, thanks to the extra $+o(1)$ in the exponent. $\mathrm{subw}(Q)$ denotes the submodular width of $Q$. . For acyclic queries, we have $\mathsf{\omega}(Q)=1$ (Yannakakis, 1981). For the triangle query, $\mathsf{\omega}(Q_{\triangle})\geq\frac{4}{3}$ modulo the 3SUM conjecture (Patrascu, 2010). In this paper, we prove several lower bounds on the IVM problems in terms of the function $\mathsf{\omega}(Q)$. These lower bounds are not conditioned on any fine-grained complexity conjecture (like the OMv conjecture (Henzinger et al., 2015), stated in Appendix B). Instead, they apply no matter what the value of $\mathsf{\omega}(Q)$ turns out to be. In that sense, they can be thought of as unconditional. [^3] Alternatively, these lower bounds can be thought of as conditioned on (symbolic) lower bounds on the static query evaluation problem.
@@
Consider a query $Q(\bm{X})=R_{1}(\bm{X}_{1})\wedge\ldots\wedge R_{k}(\bm{X}_{k})$ where $\bm{X}=\bigcup_{i\in[k]}\bm{X}_{i}$, and fresh variables $Z_{1},\ldots,Z_{k}$ that do not occur in $Q$. (In this paper, we drop head variables $\bm{X}$ and write $Q$ instead of $Q(\bm{X})$ for brevity.) Let $\Sigma_{k}$ be the set of permutations of the set $[k]$. For any permutation $\bm{\sigma}=(\sigma_{1},\ldots,\sigma_{n})\in\Sigma_{k}$, we denote by $\widehat{Q}_{\bm{\sigma}}$ the query that results from $Q$ by extending the schema of each atom $R_{\sigma_{i}}(\bm{X}_{\sigma_{i}})$ with the variables $Z_{1},\ldots,Z_{i}$:
@@ 3. Overview of Main Results
In this section, we overview our main results and discuss their implications. Recall that we use “query” to refer to a “natural join query” (or equivalently “full conjunctive query”). Also recall that $\mathsf{w}(Q)$ and $\mathsf{\omega}(Q)$ denote the fractional hypertree width of $Q$ and the smallest exponent for $\text{Eval}[Q]$, respectively, as per Definition 2.2. The notions of acyclicity, hierarchical, and Loomis-Whitney queries are reviewed in Sec. 4. Complexity measures are explained in Sec. 2. In our results, the constant enumeration delay is not amortized regardless of whether the update time is amortized. For all our IVM problems, $N$ is the number of single-tuple updates, while $|\mathcal{D}^{(\tau)}|$ is the database size at update time $\tau$.
@@ 3.1. Insert-Only Setting
For any acyclic query $Q$, $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\mathcal{O}(1)$ amortized update time.
@@
We accompany the above upper bound with a lower bound on the insertion time in terms of the lower bound $\mathsf{\omega}(Q)$ on the static evaluation of $Q$. This lower bound is not conditioned on a fine-grained complexity conjecture (like OMv). It holds no matter what the value of $\omega(Q)$ is:
Theorem 3.3.
For any query $Q$ and any constant $\gamma>0$, neither $\text{IVM}^{+}[Q]$ nor $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(N^{\mathsf{\omega}(Q)-1-\gamma})$ amortized update time.
@@ 3.2. Insert-Delete Setting
For any query $Q$, both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(N^{\mathsf{w}(\widehat{Q})-1})$ amortized update time, where $\widehat{Q}$ is the multivariate extension of $Q$ and $N$ is the number of single-tuple updates.
As a corollary, we prove a stronger upper bound where $N$ is replaced by $|\mathcal{D}^{(\tau)}|$, the current database size at update time $\tau$. Note that $|\mathcal{D}^{(\tau)}|\leq N$ and could be unboundedly smaller than $N$:
Corollary 3.5.
For any query $Q$, both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1})$ amortized update time, where $\widehat{Q}$ is the multivariate extension of $Q$, and $|\mathcal{D}^{(\tau)}|$ is the current database size at update time $\tau$.
Sec. 7.1 overviews our algorithm that meets the upper bound in Theorem 3.4. It involves intersection joins (Sec. 6). The following statements shed light on the relationship between a query $Q$ and its multivariate extension $\widehat{Q}$ (their proofs are deferred to Appendix F):
@@
Immediate corollaries of these statements are that, in the insert-delete setting, our approach needs: (1) amortized $\widetilde{\mathcal{O}}(1)$ update time for hierarchical queries; and (2) amortized $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ update time for the triangle query, which is the Loomis-Whitney query of degree 3. These update times mirror those given in prior work on $q$-hierarchical queries (Berkholz et al., 2017) and the full triangle query (Kara et al., 2020a), albeit our setting is restricted to full queries and initially empty databases and our update times are amortized and have a polylog factor.
A close analysis of the lower bound proof for non-hierarchical queries in prior work (Berkholz et al., 2017) reveals that for any non-hierarchical query $Q$ and any $\gamma>0$, there is no algorithm that solves $\text{IVM}^{\pm}[Q]$ or $\text{IVM}^{\pm}_{\delta}[Q]$ with $\widetilde{\mathcal{O}}(N^{1/2-\gamma})$ amortized update time, unless the OMv-conjecture fails (Proposition F.5 in Appendix F). Following Corollary 3.5 and the fact that $\mathsf{w}(\widehat{Q})=\frac{3}{2}$ for any Loomis-Whitney query $Q$, we conclude that both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ amortized update time, which is optimal, unless the OMv conjecture fails.
@@
We complement the upper bound on the update time in Theorem 3.4 with a lower bound in terms of the lower bound $\mathsf{\omega}(\widehat{Q}_{\bm{\sigma}})$ for the static evaluation of any component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$. This lower bound is not conditioned on any complexity conjecture. However, it only applies to the $\text{IVM}^{\pm}_{\delta}[Q]$ problem. Sec. 7.2 gives the high-level idea:
@@
Since $|\mathcal{D}^{(\tau)}|\leq N$, the above immediately implies that $\text{IVM}^{\pm}_{\delta}[Q]$ cannot be solved with amortized update time $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{\omega}(\widehat{Q}_{\bm{\sigma}})-1-\gamma})$.
4. Preliminaries
Data Model
A schema is a tuple of variables. We treat schemas and sets of variables interchangeably, assuming a fixed ordering of variables. The domain of a variable $X$ is denoted by $\text{\sf Dom}(X)$. A value tuple $\bm{t}$ over schema $\bm{X}=(X_{1},\ldots,X_{n})$ is an element from $\text{\sf Dom}(\bm{X})\stackrel{{\scriptstyle\text{def}}}{{=}}\text{\sf Dom}(X_{1})\times\cdots\times\text{\sf Dom}(X_{n})$. We denote by $\bm{t}(X)$ the $X$-value of $\bm{t}$. A relation over schema $\bm{X}$ is a finite set of value tuples over the same schema. The size $|R|$ of a relation $R$ is the number of its tuples. A database $\mathcal{D}$ is a finite set of relations. Its size $|\mathcal{D}|$ is given by the sum of the sizes of its relations.
Queries
(5) $Q(\bm{X})=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{k}(\bm{X}_{k}),$
where each $R_{i}$ is a relation symbol, each $\bm{X}_{i}$ is a tuple of variables, and $\bm{X}=\bigcup_{i\in[k]}\bm{X}_{i}$. We call each $R_{i}(\bm{X}_{i})$ an atom of $Q$. We denote by $\text{\sf at}(Q)=\{R_{i}(\bm{X}_{i})\mid i\in[k]\}$ the set of all atoms of $Q$ and by $\text{\sf at}(X)$ the set of atoms $R_{i}(\bm{X}_{i})$ with $X\in\bm{X}_{i}$. The set of all variables of $Q$ is denoted by $\text{\sf vars}(Q)=\bigcup_{i\in[k]}\bm{X}_{i}$. Since the set $\bm{X}$ is always equal to $\text{\sf vars}(Q)$ for all queries we consider in the paper, we drop $\bm{X}$ from the query notation and abbreviate $Q(\bm{X})$ with $Q$ in Eq. (5). We say that $Q$ is without self-joins if every relation symbol appears in at most one atom.
Query Classes
Acyclic queries are reviewed in Appendix B.
Single-tuple Updates and Deltas
Constant-Delay Enumeration
Width Measures
The AGM-bound and fractional hypertree width of join queries are defined as usual. Their definitions are deferred to Appendix B. If a query is a union of join queries, its AGM-bound (or fractional hypertree width) is the maximum AGM-bound (or fractional hypertree width) over its components.
@@ 5. IVM: Insert-Only Setting
In this section, we give an overview of our algorithm for $\text{IVM}^{+}[Q]$ that meets the upper bound of Theorem 3.1. We leave the details and proofs (including the extension to $\text{IVM}^{+}_{\delta}[Q]$) to Appendix C. We start with the following lemma: (Recall notation from Sec. 4.)
@@
Note that the total computation time above is the same as the AGM bound (Atserias et al., 2013) of $Q$ over the final database $\mathcal{D}^{(N)}$. In particular, even in the static setting where $\mathcal{D}^{(N)}$ is given upfront, the output size $|Q(\mathcal{D}^{(N)})|$ could be as large as $\text{\sf AGM}(Q,\mathcal{D}^{(N)})$ in the worst-case. Moreover, worst-case optimal join algorithms cannot beat this runtime (Ngo et al., 2018; Veldhuizen, 2014; Ngo, 2018). The above lemma is proved based on the query decomposition lemma (Ngo et al., 2014; Ngo, 2018), which says the following. Let $Q$ be a query and let $\bm{Y}\subseteq\text{\sf vars}(Q)$. Then, the AGM bound of $Q$ can be decomposed into a sum of AGM bounds of “residual” queries: one query $Q\Join\bm{y}$ for each tuple $\bm{y}$ over the variables $\bm{Y}$. Appendix C gives a proof of the query decomposition lemma and an example.
@@
In the above example, we only needed to do the calibration bottom-up from the leaf $Q_{1}$ to the root $Q_{2}^{\prime}$. We show in Appendix C.4 that this is not sufficient for $\text{IVM}^{+}_{\delta}[Q]$. Instead, we also need to do an extra calibration top-down. For example, suppose we have an insert of a tuple $(a,b)$ into $R$ above. In order to enumerate the new output tuples corresponding to $(a,b)$, we have to start our enumeration from $Q_{1}$ as the root. For that purpose, $Q_{1}$ also needs to be calibrated with $Q_{2}^{\prime}$. See Example C.3.

### v4

@@ 3 $+T(a_{1},c_{1})$$\{(a_{1},b_{1},c_{1})\}$$\{+Q(a_{1},b_{1},c_{1})\}$
Given a query $Q$, let $\mathsf{w}(Q)$ denote the fractional hypertree width of $Q$ (see Section 2). Let $\mathsf{\omega}(Q)$ denote the smallest exponent $\kappa$ such that $\text{Eval}[Q]$ has an algorithm with runtime $\mathcal{O}(|\mathcal{D}|^{\kappa+o(1)}+|Q(\mathcal{D})|)$ on any input database $\mathcal{D}$. Given a union of queries $Q=\bigvee_{i}Q_{i}$, we define $\mathsf{w}(Q)$ and $\mathsf{\omega}(Q)$ to be the maximum of $\mathsf{w}(Q_{i})$ and $\mathsf{\omega}(Q_{i})$, respectively, over all $i$.  [^1] The function $\omega(Q)$ is defined analogously to the constant $\omega$ which is the best exponent for matrix multiplication.
The function $\mathsf{\omega}(Q)$ is not known in general. However, we know that $\mathsf{\omega}(Q)\leq\mathsf{w}(Q)$; see e.g. (Abo Khamis et al., 2016). More strongly, $\mathsf{\omega}(Q)$ is upper bounded by the submodular width of $Q$ (Marx, 2013; Abo Khamis et al., 2017; Abo Khamis et al., 2024)222 (Abo Khamis et al., 2017; Abo Khamis et al., 2024) provide a query evaluation algorithm with runtime $\mathcal{O}(|\mathcal{D}|^{\mathrm{subw}(Q)}\cdot\text{\sf polylog}{|\mathcal{D}|}+|Q(\mathcal{D})|)=\mathcal{O}(|\mathcal{D}|^{\mathrm{subw}(Q)+o(1)}+|Q(\mathcal{D})|)$, thanks to the extra $+o(1)$ in the exponent. $\mathrm{subw}(Q)$ denotes the submodular width of $Q$. . For acyclic queries, we have $\mathsf{\omega}(Q)=1$ (Yannakakis, 1981). For the triangle query, $\mathsf{\omega}(Q_{\triangle})\geq\frac{4}{3}$ modulo the 3SUM conjecture (Patrascu, 2010). In this paper, we prove lower bounds on the IVM problems in terms of the function $\mathsf{\omega}(Q)$. These lower bounds are not conditioned on a fine-grained complexity conjecture, like the OMv conjecture (Henzinger et al., 2015) stated in Appendix B. Instead, they apply no matter what the value of $\mathsf{\omega}(Q)$ turns out to be. In that sense, they can be thought of as unconditional. However, they are relatively straightforward to prove and are mostly meant to justify the matching upper bounds.
@@
Consider a query $Q(\bm{X})=R_{1}(\bm{X}_{1})\wedge\ldots\wedge R_{k}(\bm{X}_{k})$ where $\bm{X}=\bigcup_{i\in[k]}\bm{X}_{i}$, and fresh variables $Z_{1},\ldots,Z_{k}$ that do not occur in $Q$. (In this paper, we drop head variables $\bm{X}$ and write $Q$ instead of $Q(\bm{X})$ for brevity.) Let $\Sigma_{k}$ be the set of permutations of the set $[k]$. For any permutation $\bm{\sigma}=(\sigma_{1},\ldots,\sigma_{k})\in\Sigma_{k}$, we denote by $\widehat{Q}_{\bm{\sigma}}$ the query that results from $Q$ by extending the schema of each atom $R_{\sigma_{i}}(\bm{X}_{\sigma_{i}})$ with the variables $Z_{1},\ldots,Z_{i}$:
@@ 4. Overview of Main Results
In this section, we overview our main results and discuss their implications.
@@ 4.1. Insert-Only Setting
For any query $Q$, both $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\mathcal{O}(N^{\mathsf{w}(Q)-1})$ amortized update time and non-amortized constant enumeration delay, where $N$ is the number of single-tuple inserts.
@@
For any acyclic query $Q$, $\text{IVM}^{+}[Q]$ and $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\mathcal{O}(1)$ amortized update time and non-amortized constant enumeration delay.
@@
We accompany the upper bound from Theorem 4.1 with a lower bound on the insertion time in terms of the lower bound $\mathsf{\omega}(Q)$ on the static evaluation of $Q$ (Definition 3.2). Unlike prior lower bounds (Berkholz et al., 2017; Kara et al., 2019), this lower bound is not conditioned on a fine-grained complexity conjecture, such as the OMv conjecture (Henzinger et al., 2015). However, its proof is straightforward by viewing static evaluation as a stream of inserts. It is meant to show that Theorem 4.1 is optimal up to the gap between $\mathsf{w}(Q)$ and $\mathsf{\omega}(Q)$. 333 More broadly, our lower bounds are meant to introduce a new framework to assess the optimality of algorithms in database theory, in terms of the function $\omega(Q)$. It also applies to both amortized and non-amortized update time and enumeration delay.  [^4] Note that a lower bound on the amortized time also implies the same lower bound on the worst-case time. This is because if an amortized algorithm does not exist, then a worst-case algorithm does not exist either.
Proposition 4.3.
For any query $Q$ and any constant $\gamma>0$, neither $\text{IVM}^{+}[Q]$ nor $\text{IVM}^{+}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(N^{\mathsf{\omega}(Q)-1-\gamma})$ (amortized) update time and (non-amortized) constant enumeration delay.
@@ 4.2. Insert-Delete Setting
For any query $Q$, both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1})$ amortized update time and non-amortized constant enumeration delay, where $\widehat{Q}$ is the multivariate extension of $Q$, and $|\mathcal{D}^{(\tau)}|$ is the current database size at update time $\tau$.
Sec. 7.1 overviews our algorithm that meets the upper bound in Theorem 4.4. It involves intersection joins (Sec. 6). The following statements shed light on the relationship between a query $Q$ and its multivariate extension $\widehat{Q}$ (their proofs can be found in Appendix F):
@@
Immediate corollaries of these statements are that, in the insert-delete setting, our approach needs: (1) amortized $\widetilde{\mathcal{O}}(1)$ update time for hierarchical queries; and (2) amortized $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ update time for the triangle query, which is the Loomis-Whitney query of degree 3. These update times mirror those given in prior work on $q$-hierarchical queries (Berkholz et al., 2017) and the full triangle query (Kara et al., 2020a), albeit our setting is more restricted: it considers full conjunctive queries and initially empty databases and the update times are amortized and have a polylog factor.
A close analysis of the lower bound proof for non-hierarchical queries in prior work (Berkholz et al., 2017) reveals that for any non-hierarchical query $Q$ and any $\gamma>0$, there is no algorithm that solves $\text{IVM}^{\pm}[Q]$ or $\text{IVM}^{\pm}_{\delta}[Q]$ with amortized $\widetilde{\mathcal{O}}(N^{1/2-\gamma})$ update time, unless the OMv-conjecture fails (Proposition F.5 in Appendix F). Following Theorem 4.4 and the fact that $\mathsf{w}(\widehat{Q})=\frac{3}{2}$ for any Loomis-Whitney query $Q$, we conclude that both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ amortized update time, which is optimal, unless the OMv conjecture fails.
@@
We complement the upper bound in Theorem 4.4 with a lower bound in terms of the lower bound $\mathsf{\omega}(\widehat{Q}_{\bm{\sigma}})$ for the static evaluation of any component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$. It is not conditioned of a complexity conjecture. However, it only applies to the $\text{IVM}^{\pm}_{\delta}[Q]$ problem. In particular, it is meant to show that our upper bound from Theorem 4.4 is tight for $\text{IVM}^{\pm}_{\delta}[Q]$ up to the gap between $\mathsf{w}(\widehat{Q})$ and $\mathsf{\omega}(\widehat{Q}_{\bm{\sigma}})$. It also applies to both amortized and non-amortized update time and enumeration delay. Sec. 7.2 gives the high-level idea:
@@
Let $Q$ be a query and $\widehat{Q}_{\bm{\sigma}}$ any component of its multivariate extension. For any constant $\gamma>0$, $\text{IVM}^{\pm}_{\delta}[Q]$ cannot be solved with (amortized) update time $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{\omega}(\widehat{Q}_{\bm{\sigma}})-1-\gamma})$ and (non-amortized) constant enumeration delay.
@@ 5. IVM: Insert-Only Setting
In this section, we give an overview of our algorithm for $\text{IVM}^{+}[Q]$ that meets the upper bound of Theorem 4.1. We leave the details and proofs to Appendix C. We start with the following lemma (Recall notation from Sec. 2):
@@
The total computation time above is the same as the AGM bound (Atserias et al., 2013) of $Q$ over the final database $\mathcal{D}^{(N)}$. In particular, even in the static setting where $\mathcal{D}^{(N)}$ is given upfront, the output size $|Q(\mathcal{D}^{(N)})|$ can be as large as $\text{\sf AGM}(Q,\mathcal{D}^{(N)})$ in the worst-case. Moreover, worst-case optimal join algorithms cannot beat this runtime (Ngo et al., 2018; Veldhuizen, 2014; Ngo, 2018). The above lemma is proved based on the query decomposition lemma (Ngo et al., 2014; Ngo, 2018), which says the following. Let $Q$ be a query and let $\bm{Y}\subseteq\text{\sf vars}(Q)$. Then, the AGM bound of $Q$ can be decomposed into a sum of AGM bounds of “residual” queries: one query $Q\Join\bm{y}$ for each tuple $\bm{y}$ over the variables $\bm{Y}$. Appendix C gives a proof of the query decomposition lemma and an example.
Lemma 5.2 (Query Decomposition Lemma (Ngo et al., 2014; Ngo, 2018)).
(9) $\sum_{\bm{y}\in\text{\sf Dom}(\bm{Y})}\underbrace{\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R(\bm{X})\ltimes\bm{y}|^{\lambda_{R(\bm{X})}}}_{\text{AGM-bound of $Q\ltimes\bm{y}$}}\leq\underbrace{\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R|^{\lambda_{R(\bm{X})}}}_{\text{AGM-bound of $Q$}}$
@@
In the above example, we only calibrate bottom-up from the leaf $Q_{1}$ to the root $Q_{2}^{\prime}$. However, this is not sufficient for $\text{IVM}^{+}_{\delta}[Q]$. Instead, we also need to calibrate top-down, as the following example demonstrates.
Example 5.5 (for Theorem 4.1, $\text{IVM}^{+}_{\delta}$).
Consider again the query (12) from Example 5.4. Suppose now that we want to extend our solution from Example 5.4 to the $\text{IVM}^{+}_{\delta}[Q]$ problem. In particular, suppose we have an insert of a tuple $(a,b)$ into $R$ above. In order to enumerate the new output tuples corresponding to $(a,b)$, we have to start our enumeration from $Q_{1}$ as the root. For that purpose, $Q_{1}$ also needs to be calibrated with $Q_{2}^{\prime}$. To that end, in addition to $Q_{1},P_{1}$ and $Q_{2}^{\prime}$ that are defined in Example 5.4, we also need to maintain the following relations:
$P_{1}^{\prime\prime}(B,C)$ $=Q_{2}^{\prime}(B,C,D)$